# Phase structure of (3+1)-dimensional dense two-color QCD at $T=0$ in the strong coupling limit with tensor renormalization group

### v1

@@ Acknowledgements.
Numerical calculation for the present work was carried out using the computational resources of SQUID provided by Osaka University through the HPCI System Research Project (Project ID: hp250120). We also used Yukawa-21 at Yukawa Institute Computer Facility in Kyoto University and the supercomputer Pegasus under the Multidisciplinary Cooperative Research Program of Center for Computational Sciences, University of Tsukuba. Y.S. acknowledges support from the Graduate Program on Physics for the Universe (GP-PU), Tohoku University, and from JSPS KAKENHI (Grant-in-Aid for JSPS Fellows) Grant Number 25KJ0537. SA acknowledges the support from JSPS KAKENHI Grant Number JP23K13096, and the Top Runners in Strategy of Transborder Advanced Researches (TRiSTAR) program conducted as the Strategic Professional Development Program for Young Researchers by the MEXT. This work is supported in part by Grants-in-Aid for Scientific Research from the Ministry of Education, Culture, Sports, Science and Technology (MEXT) (Nos. 24H00214, 24H00940).
@@ Appendix A Explicit form of $F$ and $R$
After performing the integration, the auxiliary Grassmann variables $\zeta,\xi$ appear in the order
$\left[\prod_{\nu}\left(\bar{\zeta}^{2}_{\nu}\chi^{a_{\nu}}\right)^{i^{\prime 2}_{\nu}}\right]\left[\prod_{\nu}\left(\bar{\zeta}^{1}_{\nu}\chi^{b_{\nu}}\right)^{i^{\prime 1}_{\nu}}\right]\left[\prod_{\nu}\left(\xi^{1}_{\nu}\chi^{1}\right)^{j^{1}_{\nu}}\right]\left[\prod_{\nu}\left(\xi^{2}_{\nu}\chi^{2}\right)^{j^{2}_{\nu}}\right]$ (24)
@@
Finally, we briefly comment on the numerical strategy for constructing the initial tensor in Eq. (2.2). Since the size of $T$ is $16^{8}$, it is impractical to implement a naive for-loop over all indices. However, careful inspection of the Kronecker delta constraints arising from the fermion and link variable integrations in Eq. (2.2) shows that they restrict nonzero contributions to only a small subset of entries, yielding a highly sparse and low-rank structure. Therefore, we performed for-loops only over the indices allowed by the Kronecker delta constraints and constructed $T$ in the form of a sparse matrix. To generate the fundamental tensors required in the ATRG method, we applied truncated SVD to $T$ using Arpack doi:10.1137/1.9780898719628 .

### v2

@@ Acknowledgements.
Numerical calculation for the present work was carried out using the computational resources of SQUID provided by Osaka University through the HPCI System Research Project (Project No. hp250120). We also used Yukawa-21 at Yukawa Institute Computer Facility in Kyoto University and the supercomputer Pegasus under the Multidisciplinary Cooperative Research Program of Center for Computational Sciences, University of Tsukuba. Y. S. acknowledges support from the Graduate Program on Physics for the Universe (GP-PU), Tohoku University, and from JSPS KAKENHI (Grant-in-Aid for JSPS Fellows) Grant No. 25KJ0537. S. A. acknowledges the support from JSPS KAKENHI Grant No. JP23K13096, and the Top Runners in Strategy of Transborder Advanced Researches (TRiSTAR) program conducted as the Strategic Professional Development Program for Young Researchers by the MEXT. This work is supported in part by Grants-in-Aid for Scientific Research from the Ministry of Education, Culture, Sports, Science and Technology (MEXT) (Grants No. 24H00214, No. 24H00940).
@@ Appendix A EXPLICIT FORM OF $F$ AND $R$
After performing the integration of link variables and staggered fields, the auxiliary Grassmann variables $\zeta,\xi$ appear in the order
$\left[\prod_{\nu}\left(\bar{\zeta}^{2}_{\nu}\right)^{i^{\prime 2}_{\nu}}\right]\left[\prod_{\nu}\left(\bar{\zeta}^{1}_{\nu}\right)^{i^{\prime 1}_{\nu}}\right]\left[\prod_{\nu}\left(\xi^{1}_{\nu}\right)^{j^{1}_{\nu}}\right]\left[\prod_{\nu}\left(\xi^{2}_{\nu}\right)^{j^{2}_{\nu}}\right]$
$\times\left[\prod_{\nu}\left(\bar{\xi}^{2}_{\nu}\right)^{j^{\prime 2}_{\nu}}\right]\left[\prod_{\nu}\left(\bar{\xi}^{1}_{\nu}\right)^{j^{\prime 1}_{\nu}}\right]\left[\prod_{\nu}\left(\zeta^{1}_{\nu}\right)^{i^{1}_{\nu}}\right]\left[\prod_{\nu}\left(\zeta^{2}_{\nu}\right)^{i^{2}_{\nu}}\right].$ (24)
@@
Finally, we briefly comment on the numerical strategy for constructing the initial tensor in Eq. (II.2). Since the size of $T$ is $16^{8}$, it is impractical to implement a naïve for-loop over all indices. However, careful inspection of the Kronecker delta constraints arising from the fermion and link variable integrations in Eq. (II.2) shows that they restrict nonzero contributions to only a small subset of entries, yielding a highly sparse and low-rank structure. Therefore, we performed for-loops only over the indices allowed by the Kronecker delta constraints and constructed $T$ in the form of a sparse matrix. To generate the fundamental tensors required in the ATRG method, we applied truncated SVD to $T$ using Arpack [46]. Once the five-leg tensors required for the subsequent ATRG calculation are constructed, they are treated as dense matrices endowed with a graded $\mathbb{Z}_{2}$ symmetry originating from the Grassmann parity throughout the coarse-graining procedure.